# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ Abstract
Incompressible insulating phases of electronic systems at partial filling of a lattice are often associated with charge ordering that breaks lattice symmetry. The resulting phases have an enlarged unit cell with an effective integer filling. Here we explore the possibility of insulating states – which we dub “Quantum Charge Liquids” (QCL) – at partial lattice filling that preserve lattice translation symmetry. Such QCL phases must necessarily either have gapped fractionally charged excitations and associated topological order or have gapless neutral excitations. We establish some general constraints on gapped fermionic QCL phases that restrict the nature of their topological order. We prove a number of results on the minimal topological order that is consistent with the lattice filling. In particular we show that at rational fillings $\nu=p/q$ with $q$ an even integer the minimal ground state degeneracy on a torus of the fermionic QCL is $4q^{2}$, 4 times larger than that of the bosonic QCL at the same filling. We comment on models and physical systems which may host fermionic QCL phases and discuss the phenomenology of these phases.
@@ I Introduction
The competition between inter-electron Coulomb repulsion and electronic kinetic energy is at the heart of many interesting phenomena in condensed matter physics. If the electrons live on a crystalline lattice, then at any fractional filling $\nu$, defined as the mean number of electrons per unit cell, metallic Fermi liquid phases will occur if the kinetic energy dominates. When the Coulomb energy dominates, and the filling $\nu$ is rational, an insulating state is expected where the electrons are localized in a spatial pattern that breaks lattice translation symmetry. These states are known as Wigner-Mott insulators. Recent experiments Regan et al. (2020); Wang et al. (2020); Xu et al. (2020); Mak and Shan (2022) on moiré transition metal dichalcogenide (TMD) materials have found evidence for Wigner-Mott insulating phases at many fractional fillings $\nu$.
@@
In this paper we focus on insulating states at fractional lattice filling. We explore the possibility of states which preserve the translation symmetry of the lattice and are therefore distinct from Wigner-Mott insulators. The well-known Lieb-Schultz-Mattis-Oshikawa-Hastings (LSMOH) theorem Lieb et al. (1961); Oshikawa (2000); Hastings (2004) tells us that such states necessarily have fractional charge excitations and associated topological order (TO), or have gapless neutral excitations. We will refer to these as ‘Quantum Charge Liquids’ (QCL). We specifically focus on the case where the intervening phase is fully gapped, and hence topological, as indicated in Fig. 1.
@@
Figure 1: A possible evolution between the Wigner-Mott insulator and a Fermi liquid at a fixed fractional lattice filling. $t$ is an electron hopping parameter that sets the scale of the bandwidth, and $U$ is a measure of the electron-electron interaction strength. We consider an intermediate phase where the charge order is lost before the insulator-metal transition happens upon increasing $t/U$. This is displayed in the cartoon, where the dimensionless charge gap is indicated by $\Delta_{c}/U$, while the dimensionless order parameter of the $\sqrt{3}\times\sqrt{3}$ charge density wave is indicated by $\rho_{\sqrt{3}\times\sqrt{3}}$ and goes to zero before the charge gap closes. General constraints tell us that this insulating phase can be fully gapped but must exhibit TO, and have excitations that carry fractional charge. For a treatment of the possibility of a continuous phase transition between the metal and the Wigner-Mott insulator without this intervening topologically ordered phase, see Refs. Xu et al. (2022); Musser et al. (2022); Musser and Senthil (2022).
@@
Though this paper is focused on the experimentally relevant case of fermionic QCLs, we note that the concept, if not the name, of QCLs has arisen before in bosonic systems. One early example is the triangular lattice quantum dimer model, which can be mapped to a hopping problem of hardcore bosons on a kagome lattice Moessner and Raman (2008). Other tractable microscopic models for bosonic fractionalized phases have also been constructed Balents et al. (2002); Senthil and Motrunich (2002); Motrunich and Senthil (2002); Motrunich (2003). More fractionalized states of bosons can be accessed theoretically either through parton constructions Senthil and Fisher (2000), or through vortex condensation from a proximate superfluid phase Balents et al. (1999); Senthil and Fisher (2000); Lannert et al. (2001); Balents et al. (2005). [^1] We note that many of these constructions are linked to closely related states Read and Sachdev (1991); Wen (1991); Savary and Balents (2016); Broholm et al. (2020) that occur as gapped Quantum Spin Liquid (QSL) phases of quantum magnets with, at least, a $\mathrm{U}(1)$ spin symmetry. One key difference between quantum magnets and the charge frustrated systems of interest in this paper is the action of time reversal symmetry. Electrical charge is even under time reversal while the spin is odd under time reversal. Thus time reversal acts differently on the generator of global $\mathrm{U}(1)$ in the two cases. The incompressible insulating states of interest in this paper where the lattice filling of particles is a generic rational fraction correspond, in the quantum magnet analog, to magnetization plateau states which, except in the special case of zero magnetization, are not time reversal invariant. Thus we can even have chiral TO as a possibility for magnetization plateau states Misguich et al. (2001) while they are not allowed at fractional filling of a time reversal invariant charge system. Apart from this conceptual distinction, the natural microscopic interactions of charge frustrated systems and quantum magnets are of course different. Finally the concept of a fermionic QCL discussed in this paper has no analog in a spin system..
@@
In contrast, fermionic QCLs have gotten surprisingly little attention despite their relevance to the aforementioned experiments in moiré TMDs. Early works Pollmann (2006); Pollmann et al. (2006, 2011) suggest that spinless fermions at half-filling on a lattice can form a $\mathbb{Z}_{2}$ TO insulating state that preserves charge conservation, lattice translation, and time reversal symmetries. Contrary to these suggestions, and to the familiar bosonic case, we will show that such a state is not allowed to exist on general grounds.
@@
For concreteness we now specialize to spinless fermions. This is relevant to the TMD materials where the spin exchange scale is very small compared to the charge exchange scale Mak and Shan (2022). The spins are thus easily polarized in a small magnetic field, or become incoherent at small temperatures between the spin and charge exchange scales. With this assumption, we present general results about the possible TOs the spinless fermions can form by extending the LSMOH constraints to the fermionic case where: $\mathrm{U}(1)$, lattice translation, and time reversal symmetries are respected. We remark that fractionalized states of fermions on a lattice in the absence of time reversal symmetry (either due to explicit or spontaneous breaking) occur in fractional quantum anomalous Hall/fractional Chern insulator states that have been observed recently Park et al. (2023); Zeng et al. (2023); Lu et al. (2024). As a warm-up we will also obtain some results on these time reversal broken fractionalized states. We will show that in general the “minimal order”, which we define below, realized by spinless fermions is larger than that of bosons.
@@
Given that LSMOH requires nontrivial TO, we can ask what the simplest or minimal TO is consistent with the lattice filling and symmetries. We propose using ground state degeneracy (GSD) on the torus to define [^2] The notion of minimal TO has appeared before in other contexts; see Ref. Sodemann et al. (2017). minimal TOs rather than quantum dimension, as done in a recent paper Jian et al. (2024). We then show that for fillings of $\nu=p/q$ per unit cell with $p,q$ relatively co-prime integers, and $q$ even, the minimal fermionic TO that does not break either translation or time reversal has a GSD on the torus of at least $4q^{2}$, larger by a factor of four than the minimal GSD of a bosonic TO at the same filling. This was discussed using dual vortex and fermionic tensor network arguments for the special case of $\nu=1/2$ in Bultinck and Cheng (2018), though it was not proven. In contrast for fillings of $\nu=p/q$ with $q$ an odd integer the minimal fermionic TO will have a GSD of at least $q^{2}$, exactly the minimal GSD of a bosonic TO at the same filling. We give a proof of these statements, and physically justify the emergence of such TOs using parton arguments. We then prove that there is only a single minimal TO consistent with: lattice translation, charge conservation, and time reversal at these fillings, so the minimal TO is unique.
@@
While our focus in this paper is on gapped QCLs, where we can establish some rigorous results, it is also interesting to contemplate the physics of gapless QCL phases. If we are interested in insulating QCLs, the charge must be fully gapped. Thus the gapless sector, if present at all, must consist entirely of electrically neutral excitations [^3] This is yet another difference with QSL states of interest in quantum magnetism. Insulating quantum spin liquids may well have gapless spin excitations. A specific example of such a gapless insulating QCL with an emergent Fermi surface of neutral fermions is described in App. B. These may well also appear as intermediate phases between the Wigner-Mott insulator and a Fermi liquid. Finally we could also consider metallic QCLs. The Fermi liquid itself could be thought of as a familiar example; non-fermi liquid metallic states that preserve the symmetries and filling constraints will surely qualify as metallic QCLs.
@@
Where might we expect to see fermionic QCLs? In the better understood bosonic systems, fractionalized insulators typically occur in the ‘intermediate’ correlation limit where the kinetic energy leads to ring exchange processes. By analogy, and as motivated above, two dimensional moiré TMD materials evolving between Wigner-Mott insulators and Fermi liquids might be an interesting platform to observe fractionalized QCL insulators. There has also been a recent suggestion Mao et al. (2023) that twisted bilayer graphene at $1/3$ filling might realize a fractionalized phase. Our results on the allowed TO phases at fractional filling should inform further explorations of this suggestion. Finally we note that the physics of charge frustration appears in a few traditional solid state systems as well, e.g., in triangular lattice quarter-filled organic materials Hotta (2012). Though not as tunable as moire materials, QCL phases might occur in some such systems as well.
@@
The outline of the rest of the paper is as follows. In Sec. II we give a brief review of the mathematical theory of TO and define different classes of “minimal” orders. In Sec. III we find the minimal orders at a fixed fractional filling of a lattice $\nu=1/q$ with and without the presence of time reversal symmetry. With time reversal symmetry we will show that there is a unique minimal TO, $\mathbb{Z}_{2q}$ ($\mathbb{Z}_{q}$) gauge theory for $q$ even (odd). Finally in Sec. IV we discuss areas of further study in three related directions: 1) finding microscopic models that realize gapped fermionic QCLs, 2) describing the phenomenology of such QCLs, including their response to doping, and 3) other scenarios where the concept of minimal order could prove clarifying, as well as theoretical questions raised by the concept of minimal order.
@@ II Definition of “minimal” symmetry enriched topological order
We first give a brief review of the theory of TO to establish some common definitions. For a more comprehensive review, see Kitaev (2006) and Bonderson (2007). A TO $\mathcal{C}$ is characterized by its anyon types and the chiral central charge $c_{-}$. We will focus on the anyons, since $c_{-}$ must be 0 in time reversal invariant systems, which is the main subject of the work. Two anyons $a,b\in\mathcal{C}$ can be fused, resulting in a direct sum of other anyons $a\times b=\sum_{c}N^{c}_{ab}c$, where $N^{c}_{ab}\geq 0$ are integers known as the fusion multiplicity. The fusion rules determine the quantum dimensions $d_{a}\geq 1$ for the anyons, through the equation $d_{a}d_{b}=\sum_{c}N_{ab}^{c}d_{c}$. If $d_{a}=1$ then $a$’s fusion outcome with any other anyon is unique and it is thus an Abelian anyon. One important Abelian anyon is the unique “identity” anyon $1$, which corresponds to a local bosonic excitation. It has the property that $1\times a=a$ for all $a\in\mathcal{C}$. The set of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ forms an Abelian group where the group multiplication is defined by fusion and the identity element is given by $1$. If $\mathcal{A}=\mathcal{C}$ then we refer to $\mathcal{C}$ as an Abelian TO. The total quantum dimension $\mathcal{D}_{\mathcal{C}}$ of $\mathcal{C}$ is defined as
@@
In addition to the fusion rules for anyons in $\mathcal{C}$ a TO also needs to specify the exchange and braiding statistics of the anyons. The exchange statistics of an anyon $a$ is denoted by $e^{i\theta_{a}}$, where $\theta_{a}$ is referred to as the topological spin of $a$. The effect of braiding an anyon $a$ around $b$ is captured by the monodromy $M_{ab}$. If either $a$ or $b$ is an Abelian anyon then $M_{ab}=e^{i\theta_{a,b}}$, where $\theta_{a,b}$ is referred to as the braiding phase. Restricted to $\mathcal{A}\subseteq\mathcal{C}$ the function $\theta_{a,b}$ is a symmetric bilinear form on $\mathcal{A}$, i.e. a map from $\mathcal{A}\times\mathcal{A}$ to $\mathbb{R}/2\pi\mathbb{Z}=\mathbb{S}^{1}$ that satisfies $\theta_{a,b}=\theta_{b,a}$ and $\theta_{a\times b,c}=\theta_{a,c}+\theta_{b,c}$. The identity anyon is a boson which braids trivially with all other anyons, and thus $\theta_{1}=0$ and $\theta_{1,a}=0$ for all $a\in\mathcal{C}$. In a bosonic TO, $1$ is the only anyon type with this property. In addition a fermionic TO will have an Abelian anyon $c\in\mathcal{C}$ corresponding to a local fermion, i.e. the electron. This anyon obeys $c\times c=1$, $\theta_{c}=\pi$, and $\theta_{a,c}=0$ for all $a\in\mathcal{C}$. In a fermionic TO, $1$ and $c$ are the only anyons which braid trivially with everything else. We note that physically the electron is a local particle, and as such is sometimes not counted as an anyon in $\mathcal{C}$. However, keeping $c\in\mathcal{C}$ will prove to be a useful bookkeeping device for us, so we will adopt this convention throughout. With this convention, a trivial fermionic insulator is described by the theory $\{1,c\}$.
@@
We further note that a bosonic TO $\mathcal{C}_{b}$ can be trivially turned into the fermionic TO, $\mathcal{C}_{b}\boxtimes\{1,c\}$, where $\boxtimes$ means stacking of two anyon theories. Physically, it means that the TO $\mathcal{C}_{b}$ is made of bosonic degrees of freedom in a fermionic system. However, not all fermionic TOs can be written in this form [^4] A familiar example of a fermionic TO that is not equivalent to a bosonic TO stacked with $\{1,c\}$ is the Moore-Read Pfaffian state., although Abelian ones can (see Theorem C.27 in App. C).
@@
Consistency of the fusion and braiding rules establishes constraints on the possible TOs. For a comprehensive review, see Appendix E of Kitaev (2006). If, in addition, the TO is required to respect a symmetry $G$, this will impose further constraints, ruling out some TOs Barkeshli et al. (2019). Those that are consistent with a symmetry action of $G$ are referred to as symmetry-entriched TOs (SETs). A natural question is what the “least complicated”, or minimal, SETs are given a symmetry $G$ and certain physical properties. Surprisingly, despite the well developed study of SETs, the question of minimality is a relatively recent one. Answering it first requires one to define what is meant by minimal. A recent paper Jian et al. (2024) endeavored to do this by defining a minimal SET to be one which has the smallest quantum dimension possible, given that it must respect $G$. We propose another definition of minimal. We say that a SET is minimal if it has the smallest GSD on a torus, given that it must respect $G$. We prefer our definition for two reasons. The first is that a lower bound on GSD provides a more useful tool for numerics such as exact diagonalization, where GSD on a torus is easily assessed. The second reason to prefer our definition is that it is more powerful when proving results about minimal Abelian TOs. We make this more precise below.
@@
Figure 2: A cartoon picture of the minimal order classes. The top black box is the set of SETs consistent with the symmetry $G$, SETG. Contained within SETG, illustrated by bags, are the SETs, SET1 and SET2 along with other possibilities. Each of these SETs contains anyons, illustrated by the colorful balls. The radius of each anyon is a proxy for its quantum dimension, with the smaller balls therefore being Abelian and the larger light blue ball being non-Abelian. The bottom left illustrates the minimal class $mG_{\mathcal{D}}$, containing all SETs in SETG with the smallest quantum dimension. This will correspond to the total “weight” of the balls in the bag, so SET2 is excluded because of its large non-Abelian anyon. The bottom right illustrates the minimal class $mG_{\rm GSD}$, containing all SETs with the smallest number of anyons. The large number of anyons in SET1 thus rules out its inclusion. This illustration makes clear that $mG_{\mathcal{D}}$ will tend to be biased in favor of Abelian theories, while the reverse is true of $mG_{\rm GSD}$.
Let SETG be the set of all SETs respecting $G$ and have the associated physical properties. Then define:
$mG_{\mathcal{D}}=\{\mathcal{C}\in{\rm SET}_{G}\mid\mathcal{D}_{\mathcal{C}}\leq\mathcal{D}_{\mathcal{C}^{\prime}}\ \forall\mathcal{C}^{\prime}\in{\rm SET}_{G}\},$ (2)

### v2

@@ Abstract
Incompressible insulating phases of electronic systems at partial filling of a lattice are often associated with charge ordering that breaks lattice symmetry. The resulting phases have an enlarged unit cell with an effective integer filling. Here we explore the possibility of insulating states—which we dub “quantum charge liquids” (QCL)—at partial lattice filling that preserve lattice translation symmetry. Such QCL phases must necessarily either have gapped fractionally charged excitations and associated topological order or have gapless neutral excitations. We establish some general constraints on gapped fermionic QCL phases that restrict the nature of their topological order. We prove a number of results on the minimal topological order that is consistent with the lattice filling. In particular we show that, at rational fillings $\nu=p/q$ with $q$ an even integer, the minimal ground-state degeneracy on a torus of the fermionic QCL is $4q^{2}$, four times larger than that of the bosonic QCL at the same filling. We comment on models and physical systems which may host fermionic QCL phases and discuss the phenomenology of these phases.
@@ I Introduction
The competition between interelectron Coulomb repulsion and electronic kinetic energy is at the heart of many interesting phenomena in condensed-matter physics. If the electrons live on a crystalline lattice, then at any fractional filling $\nu$, defined as the mean number of electrons per unit cell, metallic Fermi-liquid phases will occur if the kinetic energy dominates. When the Coulomb energy dominates, and the filling $\nu$ is rational, an insulating state is expected where the electrons are localized in a spatial pattern that breaks lattice translation symmetry. These states are known as Wigner-Mott insulators. Recent experiments Regan et al. (2020); Wang et al. (2020); Xu et al. (2020); Mak and Shan (2022) on moiré transition metal dichalcogenide (TMD) materials have found evidence for Wigner-Mott insulating phases at many fractional fillings $\nu$.
@@
Figure 1: A possible evolution between the Wigner-Mott insulator and a Fermi liquid at a fixed fractional lattice filling. $t$ sets the scale of the bandwidth, and $U$ is a measure of the electron-electron interaction strength. We consider an intermediate phase where the charge order is lost before the insulator-metal transition. [^2] In the extended Hubbard model $t/U\approx 0.1$ in numerical studies of the insulator-metal transition Zhou et al. (2024). This is displayed in the cartoon, where the dimensionless charge gap is $\Delta_{c}/U$ and the dimensionless charge density wave order parameter is $\rho_{\sqrt{3}\times\sqrt{3}}$. It goes to zero before the charge gap closes. General constraints tell us that this insulating phase can be fully gapped but must exhibit TO, and have excitations that carry fractional charge. For a treatment of the possibility of a continuous phase transition between the metal and the Wigner-Mott insulator without this intervening topologically ordered phase, see Refs. Xu et al. (2022); Musser et al. (2022); Musser and Senthil (2022).
In this paper we focus on insulating states at fractional lattice filling. We explore the possibility of states which preserve the translation symmetry of the lattice and are therefore distinct from Wigner-Mott insulators. The well-known Lieb-Schultz-Mattis-Oshikawa-Hastings (LSMOH) theorem Lieb et al. (1961); Oshikawa (2000); Hastings (2004) tells us that such states necessarily have fractional-charge excitations and associated topological order (TO), or have gapless neutral excitations. We will refer to these as “quantum charge liquids” (QCLs). We specifically focus on the case where the intervening phase is fully gapped, and hence topological, as indicated in Fig. 2.
@@
Although this paper is focused on the experimentally relevant case of fermionic QCLs, we note that the concept, if not the name, of QCLs has arisen before in bosonic systems. One early example is the triangular lattice quantum dimer model, which can be mapped to a hopping problem of hardcore bosons on a kagome lattice Moessner and Raman (2008). Other tractable microscopic models for bosonic fractionalized phases have also been constructed Balents et al. (2002); Senthil and Motrunich (2002); Motrunich and Senthil (2002); Motrunich (2003). More fractionalized states of bosons can be accessed theoretically either through parton constructions Senthil and Fisher (2000), or through vortex condensation from a proximate superfluid phase Balents et al. (1999); Senthil and Fisher (2000); Lannert et al. (2001); Balents et al. (2005). [^3] We note that many of these constructions are linked to closely related states Read and Sachdev (1991); Wen (1991); Savary and Balents (2016); Broholm et al. (2020) that occur as gapped quantum spin liquid (QSL) phases of quantum magnets with, at least, a $\mathrm{U}(1)$ spin symmetry Zou et al. (2018). One key difference between quantum magnets and the charge-frustrated systems of interest in this paper is the action of time-reversal symmetry. Electrical charge is even under time reversal while the spin is odd under time reversal. Thus, time reversal acts differently on the generator of global $\mathrm{U}(1)$ in the two cases. The incompressible insulating states of interest in this paper where the lattice filling of particles is a generic rational fraction correspond, in the quantum magnet analog, to magnetization plateau states which, except in the special case of zero magnetization, are not time-reversal invariant. Thus we can even have chiral TO as a possibility for magnetization plateau states Misguich et al. (2001) while they are not allowed at fractional filling of a time-reversal invariant charge system. Apart from this conceptual distinction, the natural microscopic interactions of charge-frustrated systems and quantum magnets are of course different. Finally, the concept of a fermionic QCL discussed in this paper has no analog in a spin system.
@@
In contrast, fermionic QCLs have gotten surprisingly little attention despite their relevance to the aforementioned experiments in moiré TMDs. Early works Pollmann (2006); Pollmann et al. (2006, 2011) suggest that spinless fermions at $\nu=1/2$ on a lattice can form a $\mathbb{Z}_{2}$ TO insulating state that preserves charge conservation, lattice translation, and time-reversal symmetries. Contrary to these suggestions, and to the familiar bosonic case, we show that such a state is not allowed to exist on general grounds. Instead a filling $\nu=1$ is needed to observe $\mathbb{Z}_{2}$ TO in this system, as recently found in exactly solvable models Han and Kivelson (2023); Cai et al. (2024).
@@
For concreteness we now specialize to spinless fermions. This is relevant to the TMD materials where the spin-exchange scale is very small compared with the charge-exchange scale Mak and Shan (2022). The spins are thus easily polarized in a small magnetic field, or become incoherent at small temperatures between the spin- and charge-exchange scales. With this assumption, we present general results about the possible TOs the spinless fermions can form by extending the LSMOH constraints to the fermionic case where: $\mathrm{U}(1)$, lattice translation, and time-reversal symmetries are respected. We remark that fractionalized states of fermions on a lattice in the absence of time-reversal symmetry (either due to explicit or spontaneous breaking) occur in fractional quantum anomalous Hall and fractional Chern insulator states that have been observed recently Park et al. (2023); Zeng et al. (2023); Lu et al. (2024). As a warm-up we also obtain some results on these time-reversal-broken fractionalized states. We show that, in general, the “minimal order,” which we define below, realized by spinless fermions is larger than that of bosons.
@@
Given that LSMOH requires nontrivial TO, we can ask what the simplest or minimal TO is consistent with the lattice filling and symmetries. We propose using ground-state degeneracy (GSD) on the torus to define [^4] The notion of minimal TO has appeared before in other contexts; see Ref. Sodemann et al. (2017). minimal TOs rather than quantum dimension, as done in a recent paper Jian et al. (2024). We then show that, for fillings of $\nu=p/q$ per unit cell with $p,q$ relatively coprime integers, and $q$ even, the minimal fermionic TO that does not break either translation or time reversal has a GSD on the torus of at least $4q^{2}$, larger by a factor of four than the minimal GSD of a bosonic TO at the same filling. This was discussed using dual vortex and fermionic tensor network arguments for the special case of $\nu=1/2$ in Ref. Bultinck and Cheng (2018), although it was not proven. In contrast, for fillings of $\nu=p/q$ with $q$ an odd integer, the minimal fermionic TO will have a GSD of at least $q^{2}$, exactly the minimal GSD of a bosonic TO at the same filling. We give a proof of these statements, and physically justify the emergence of such TOs using parton arguments. We then prove that there is only a single minimal TO consistent with: lattice translation, charge conservation, and time reversal at these fillings, so the minimal TO is unique.
@@
While our focus in this paper is on gapped QCLs, where we can establish some rigorous results, it is also interesting to contemplate the physics of gapless QCL phases. If we are interested in insulating QCLs, the charge must be fully gapped. Thus the gapless sector, if present at all, must consist entirely of electrically neutral excitations. [^5] This is yet another difference with QSL states of interest in quantum magnetism. Insulating quantum spin liquids may well have gapless spin excitations. A specific example of such a gapless insulating QCL with an emergent Fermi surface of neutral fermions is described in Appendix B. These may well also appear as intermediate phases between the Wigner-Mott insulator and a Fermi liquid. Finally we could also consider metallic QCLs. The Fermi liquid itself could be thought of as a familiar example; non-Fermi-liquid metallic states that preserve the symmetries and filling constraints will surely qualify as metallic QCLs.
@@
Where might we expect to see fermionic QCLs? In the better understood bosonic systems, fractionalized insulators typically occur in the “intermediate” correlation limit where the kinetic energy leads to ring exchange processes. By analogy, and as motivated above, two-dimensional moiré TMD materials evolving between Wigner-Mott insulators and Fermi liquids might be an interesting platform to observe fractionalized QCL insulators. There has also been a recent suggestion Mao et al. (2023) that twisted bilayer graphene at $1/3$ filling might realize a fractionalized phase. Our results on the allowed TO phases at fractional filling should inform further explorations of this suggestion. Finally, we note that the physics of charge frustration appears in a few traditional solid-state systems as well, e.g., in triangular lattice quarter-filled organic materials Hotta (2012). Although not as tunable as moiré materials, QCL phases might occur in some such systems as well.
@@
The outline of the rest of the paper is as follows: In Sec. II we give a brief review of the mathematical theory of TO and define different classes of “minimal” orders. In Sec. III we find the minimal orders at a fixed fractional filling of a lattice $\nu=1/q$ with and without the presence of time-reversal symmetry. With time-reversal symmetry we show that there is a unique minimal TO, $\mathbb{Z}_{2q}$ ($\mathbb{Z}_{q}$) gauge theory for $q$ even (odd). Finally in Sec. IV we discuss areas of further study in three related directions: (1) finding microscopic models that realize gapped fermionic QCLs, (2) describing the phenomenology of such QCLs, including their response to doping, and (3) other scenarios where the concept of minimal order could prove clarifying, as well as theoretical questions raised by the concept of minimal order.
@@ II Definition of “minimal” symmetry enriched topological order
We first give a brief review of the theory of TO to establish some common definitions. For a more comprehensive review, see Refs. Kitaev (2006) and Bonderson (2007). A TO $\mathcal{C}$ is characterized by its anyon types and the chiral central charge $c_{-}$. We focus on the anyons, since $c_{-}$ must be 0 in time reversal invariant systems, which is the main subject of the work. Two anyons $a,b\in\mathcal{C}$ can be fused, resulting in a direct sum of other anyons $a\times b=\sum_{c}N^{c}_{ab}c$, where $N^{c}_{ab}\geq 0$ are integers known as the fusion multiplicity. The fusion rules determine the quantum dimensions $d_{a}\geq 1$ for the anyons, through the equation $d_{a}d_{b}=\sum_{c}N_{ab}^{c}d_{c}$. If $d_{a}=1$ then $a$’s fusion outcome with any other anyon is unique and it is thus an Abelian anyon. One important Abelian anyon is the unique “identity” anyon $1$, which corresponds to a local bosonic excitation. It has the property that $1\times a=a$ for all $a\in\mathcal{C}$. The set of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ forms an Abelian group where the group multiplication is defined by fusion and the identity element is given by $1$. If $\mathcal{A}=\mathcal{C}$ then we refer to $\mathcal{C}$ as an Abelian TO. The total quantum dimension $\mathcal{D}_{\mathcal{C}}$ of $\mathcal{C}$ is defined as
@@
In addition to the fusion rules for anyons in $\mathcal{C}$ a TO also needs to specify the exchange and braiding statistics of the anyons. The exchange statistics of an anyon $a$ is denoted by $e^{i\theta_{a}}$, where $\theta_{a}$ is referred to as the topological spin of $a$. The effect of braiding an anyon $a$ around $b$ is captured by the monodromy $M_{ab}$. If either $a$ or $b$ is an Abelian anyon then $M_{ab}=e^{i\theta_{a,b}}$, where $\theta_{a,b}$ is referred to as the braiding phase. Restricted to $\mathcal{A}\subseteq\mathcal{C}$ the function $\theta_{a,b}$ is a symmetric bilinear form on $\mathcal{A}$ i.e., a map from $\mathcal{A}\times\mathcal{A}$ to $\mathbb{R}/2\pi\mathbb{Z}=\mathbb{S}^{1}$ that satisfies $\theta_{a,b}=\theta_{b,a}$ and $\theta_{a\times b,c}=\theta_{a,c}+\theta_{b,c}$. Moreover, $\theta_{a,a}=2\theta_{a}$ i.e., the braiding phase of an anyon with itself is twice its topological spin. The identity anyon is a boson which braids trivially with all other anyons, and thus $\theta_{1}=0$ and $\theta_{1,a}=0$ for all $a\in\mathcal{C}$. In a bosonic TO, $1$ is the only anyon type with this property. A fermionic TO will have an Abelian anyon $c\in\mathcal{C}$ corresponding to a local fermion i.e., the electron. This anyon obeys $c\times c=1$, $\theta_{c}=\pi$, and $\theta_{a,c}=0$ for all $a\in\mathcal{C}$. In a fermionic TO, $1$ and $c$ are the only anyons which braid trivially with everything else. We note that physically the electron is a local particle, and as such is sometimes not counted as an anyon in $\mathcal{C}$. However, keeping $c\in\mathcal{C}$ will prove to be a useful bookkeeping device for us, so we adopt this convention throughout. With this convention, a trivial fermionic insulator [^6] Here we refer to fermionic invertible states as trivial because they have no nontrivial anyons. is described by the theory $\{1,c\}$.
@@
We further note that a bosonic TO $\mathcal{C}_{b}$ can be trivially turned into the fermionic TO, $\mathcal{C}_{b}\boxtimes\{1,c\}$, where $\boxtimes$ means stacking of two anyon theories. Physically, it means that the TO $\mathcal{C}_{b}$ is made of bosonic degrees of freedom in a fermionic system. However, not all fermionic TOs can be written in this form, [^7] A familiar example of a fermionic TO that is not equivalent to a bosonic TO stacked with $\{1,c\}$ is the Moore-Read Pfaffian state. although Abelian ones can (see Theorem C.27 in Appendix C).
Consistency of the fusion and braiding rules establishes constraints on the possible TOs. For a comprehensive review, see Appendix E of Ref. Kitaev (2006). If, in addition, the TO is required to respect a symmetry $G$, this will impose further constraints, ruling out some symmetry actions and enrichments of G on the TO Barkeshli et al. (2019); Bulmash and Barkeshli (2022a, b); Aasen et al. (2022). Those that are consistent with a symmetry action of $G$ are referred to as symmetry-enriched TOs (SETs).
@@
In addition to a symmetry $G$ a system may have a “fractional topological response” Cheng et al. (2023) that requires a nontrivial TO. The example explored in this paper is fractional filling of a lattice, $\nu$, which requires nontrivial TO by the LSMOH theorem Lieb et al. (1961); Oshikawa (2000); Hastings (2004). Another familiar example is the fractional quantum Hall effect Cheng et al. (2025). A natural question is what the “least complicated,” or minimal, SETs are given a symmetry $G$ and a fractional topological response $\nu$. Surprisingly, despite the well-developed study of SETs, the question of minimality is a relatively recent one. Answering it first requires one to define what is meant by minimal. A recent paper Jian et al. (2024) endeavored to do this by defining a minimal SET to be one which has the smallest quantum dimension possible, given that it must respect $G$ and have the appropriate fractional topological response. We propose another definition of minimal. We say that a SET is minimal if it has the smallest GSD on a torus, given that it must respect $G$ and $\nu$. We prefer our definition for two reasons. The first is that a lower bound on GSD provides a more useful tool for numerics such as exact diagonalization, where GSD on a torus is easily assessed. [^8] Note that in electronic systems the GSD may have size dependence Wen (2003); Rao and Sodemann (2021); when we discuss GSD of these systems we refer to those with an even number of electrons. The second reason to prefer our definition is that it is more powerful when proving results about minimal Abelian SETs. We make this more precise below.
@@
Figure 2: A cartoon picture of the minimal order classes. The top black box is the set of SETs consistent with the symmetry $G$ and filling fraction $\nu$, SETG,ν. Contained within SETG,ν, illustrated by bags, are the SETs, SET1 and SET2 along with other possibilities. Each of these SETs contains anyons, illustrated by the colorful balls. The radius of each anyon is a proxy for its quantum dimension, with the smaller balls therefore being Abelian and the larger light blue ball being non-Abelian. The bottom-left illustrates the minimal class $mG^{(\nu)}_{\mathcal{D}}$, containing all SETs in SETG,ν with the smallest quantum dimension. This will correspond to the total “weight” of the balls in the bag, so SET2 is excluded because of its large non-Abelian anyon. The bottom-right illustrates the minimal class $mG^{(\nu)}_{\rm GSD}$, containing all SETs with the smallest number of anyons. The large number of anyons in SET1 thus rules out its inclusion. This illustration makes it clear that $mG^{(\nu)}_{\mathcal{D}}$ will tend to be biased in favor of Abelian theories, while the reverse is true of $mG^{(\nu)}_{\rm GSD}$.
Let SETG,ν be the set of all SETs respecting $G$ and $\nu$. Then define:
$mG^{(\nu)}_{\mathcal{D}}=\{\mathcal{C}\in{\rm SET}_{G,\nu}\mid\mathcal{D}_{\mathcal{C}}\leq\mathcal{D}_{\mathcal{C}^{\prime}}\ \forall\mathcal{C}^{\prime}\in{\rm SET}_{G,\nu}\},$ (2)